Calculate the △G° at 298 K for the following reaction. Show your work.


Fe2O3 (s) + 13CO (g) = 2Fe(CO)5(g) + 3CO2(g)


-824.2 -110.5 -733.8 -393.5


△H° (kJ/mol)


87.4 197.6 445.2 213.6


△S° (J/mol x K)

1
Expert's answer
2022-05-04T14:28:04-0400

Solution:

Balanced chemical equation:

Fe2O3(s) + 13CO(g) → 2Fe(CO)5(g) + 3CO2(g)


The standard enthalpy of reaction is then given by:

ΔHrxn = ∑νΔHf(products) − ∑νΔHf(reactants)

Therefore,

ΔHrxn = [3×ΔHf(CO2, g) + 2×ΔHf(Fe(CO)5, g)] − [13×ΔHf(CO, g) + ΔHf(Fe2O3, s)]

ΔHrxn = [3×(−393.5) + 2×(−733.8)] − [13×(−110.5) + (−824.2)] = −387.4

ΔHrxn = −387.4 kJ mol−1


The standard entropy of reaction is then given by:

ΔSrxn = ∑νSf(products) − ∑νSf(reactants)

Therefore,

ΔSrxn = [3×Sf(CO2, g) + 2×Sf(Fe(CO)5, g)] − [13×Sf(CO, g) + Sf(Fe2O3, s)]

ΔSrxn = [3×213.6 + 2×445.2] − [13×197.6 + 87.4] = −1125

ΔSrxn = −1125 J mol−1 K−1


ΔG = ΔH − TΔS

Therefore,

ΔGrxn = (−387400 J mol−1) − (298 K) × ( −1125 J mol−1 K−1) = −52150 J mol−1 = −52.15 kJ mol−1

ΔGrxn = −52.15 kJ mol−1


Answer: ΔGrxn = −52.15 kJ mol−1